OpenStudy (anonymous):

-3x^2+9x+7 for x = -3

OpenStudy (anonymous):

so -3 to the 2nd power plus 9x3plus 7

OpenStudy (anonymous):

36

Directrix (directrix):

-3x^2+9x+7 for x = -3 -------------------- -3* (-3)^2 + 9*(-3) + 7 = -3 * 9 - 27 + 7 which does NOT sum to 36 as stated above. @fareastmovement13